# Metabolic light absorption, scattering and emission (MetaLASE) microscopy

### v1

@@
Brendon S. Restall,${}^{1^{\dagger}}$ Nathaniel J.M. Haven,${}^{1^{\dagger}}$ Matthew T. Martell, ${}^{1^{\dagger}}$ Brendyn D. Cikaluk, ${}^{1}$ Saymon Tejay,${}^{3}$ Benjamin A. Adam,${}^{2}$ Gopinath Stendra ,${}^{3}$ Xingyu Li,${}^{1}$ Roger J. Zemp,${}^{1^{\ast}\\ }$
${}^{1}$epartment of Electrical and Computer Engineering, University of Alberta, 116 Street & 85 Avenue, Edmonton, Alberta T6G 2R3, Canada
${}^{2}$epartment of Laboratory Medicine and Pathology, University of Alberta, 8440 - 112 Street, Edmonton, Alberta T6G 2B7, Canada
${}^{3}$epartment of Medicine, Faculty of Medicine & Dentistry, University of Alberta, Edmonton, Alberta, Canada
${}^{\ast}$o whom correspondence should be addressed; E-mail: rzemp@ualberta.ca
@@ Discussion
[…]
@@
[…]

### v2

@@
Brendon S. Restall,${}^{1^{\dagger}}$ Nathaniel J.M. Haven,${}^{1^{\dagger}}$ Matthew T. Martell, ${}^{1^{\dagger}}$ Brendyn D. Cikaluk, 1 Saymon Tejay,3 Benjamin A. Adam,2 Gopinath Stendra ,3 Xingyu Li,1 Roger J. Zemp,${}^{1^{\ast}\\ }$
1Department of Electrical and Computer Engineering, University of Alberta, 116 Street & 85 Avenue, Edmonton, Alberta T6G 2R3, Canada
2Department of Laboratory Medicine and Pathology, University of Alberta, 8440 - 112 Street, Edmonton, Alberta T6G 2B7, Canada
3Department of Medicine, Faculty of Medicine & Dentistry, University of Alberta, Edmonton, Alberta, Canada
To whom correspondence should be addressed; E-mail: rzemp@ualberta.ca
@@ Discussion
To render data acquired from data acquisition channels as an image, the scan trajectory was reconstructed from AquadB encoder data using custom C++ OpenMP software for parallel read-in of encoder and optical signals, resolving stage positions from encoder state changes (fast axis) and sample timestamps (slow axis), and associating optical signal data with each position on the sinusoidal trajectory. Using Delaunay triangulation-based scattered data interpolation in a CPU-parallelized MATLAB script, this scan data was rendered onto a pixelated Cartesian grid using natural neighbor interpolation to form image arrays for each data channel. Total image formation required approximately 2.7 minutes per mm2 using a 16-thread processor (i9-9900k, Intel) with 128 GB of RAM, with considerable opportunity for speedup of the interpolation component using GPU-based acceleration, and the scan trajectory reconstruction component using hardware-based (e.g. FPGA) implementations.
@@
A549 cells were purchased and validated from ATCC (CCL-185), and cells were free from mycoplasma. Cells were cultured at 37 degrees in 9% CO2 in DMEM (Gibco #11995-073; Glucose [25 mM], Glutamine [4 mM]) supplemented with 10% FBS (Sigma-Aldrich #F1051) and 1% penstrep (Gibco #15240062). For glutamine starvation experiments, cells were grown up to 60% confluency in complete media, washed once with PBS then cultured in complete media or serum- and glutamine-free DMEM media (Gibco #A1443001) for 24 and 48 hrs on glass slides prior to imaging.
@@ Acknowledgements
Table 1: Current metabolic and virtual histology modalities along with their resolution, scan speed, and contrast mechanism. Unlike many approaches in the literature, our virtual histology approach is label free and unlike many virtual H&E microscopy implementations, our MetaLASE approach offers metabolic information. Note that specifications represent reported system implementations, but not necessarily fundamental limitations of each technology. 2P: two-photon; 3P: three-photon; SHG: second harmonic generation; THG: third harmonic generation; NLM: nonlinear microscopy; CFM: confocal fluorescence microscopy; SIM: structured illumination microscopy; FLIM: fluorescence lifetime imaging microscopy; TA-PARS: total absorption PARS; MUSE: microscopy with ultraviolet surface excitation; LSM: lightsheet microscopy; SRS: stimulated Raman scattering; PAM: photoacoustic microscopy; MS-DUV: multispectral deep UV; RCM: reflectance confocal microscopy.